OpenStudy (anonymous):

Limit ( t^3 + t^2 - 5t +3 / t^3 - 3t + 2) x tends to -1

OpenStudy (anonymous):

you mean t tends to -1?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

tell me tano plzzzz

OpenStudy (anonymous):

did you try replacing t for -1?

OpenStudy (anonymous):

yeah but it got infinity .....

OpenStudy (anonymous):

i knw its answer is 4/3 but dont knw how to solve

OpenStudy (anonymous):

the answer is 2

OpenStudy (anonymous):

if you replace t for -1 , you get 8/4 \[=2\]

OpenStudy (anonymous):

\[{(-1)^3 +(-1)^2-5(-1)+ 3\over (-1)^3-3(-1)+2} = {-1 + 1 + 5 + 3\over -1 + 3 + 2} = {8\over 4} = 2\]

OpenStudy (anonymous):

wrong dear

OpenStudy (anonymous):

(-1)^2 = 1 not -1 :(

OpenStudy (anonymous):

Pretty sure that's what I wrote..

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

polpak srry the limit was t tends to 1

OpenStudy (anonymous):

limit is 1 thn tell me solution

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

Oh.. You should say what you mean then. ;p You need to take the derivative of the top and bottom and solve for that limit. By l'Hopital's rule it will be the same.

OpenStudy (anonymous):

ok but its computing limit by simple method not by L hopital rule is there is any other way ??

OpenStudy (anonymous):

Oh, sure you can do that too.

OpenStudy (anonymous):

hmm tell me?

OpenStudy (anonymous):

w8ing

OpenStudy (anonymous):

Wait no. blah

OpenStudy (anonymous):

wrong denominator

OpenStudy (anonymous):

You have to use l'Hopital.

OpenStudy (anonymous):

:( ok

OpenStudy (anonymous):

Yeah, if it were going to 0 you would be fine with just dividing.

OpenStudy (anonymous):

But it's not, so you cannot ;p

OpenStudy (anonymous):

i was tried all this step but .... i thng L hopital rule is only way

OpenStudy (anonymous):

thnx buddy 4 ur time